OpenStudy (jhannybean):

Calculus 3: Part of the surface \(y=4x+z^2\) that lies between planes \(x=0~, ~x=1~,~z=0~,~z=1\)

OpenStudy (anonymous):

Lets see it

OpenStudy (jhannybean):

I have parametricized my equations as : \( x=x ~ , ~ y=4x+z^2 ~ , ~ z=z\)

OpenStudy (anonymous):

Yes that is correct

OpenStudy (jhannybean):

That isn't what I am looking for @Pico33

OpenStudy (jhannybean):

No. I am finding the surface Area. \[A = \int\int_D \sqrt{1+(f_x)^2 +(f_y)^2}dA\]

OpenStudy (anonymous):

Ahh ok

OpenStudy (anonymous):

Well, you've pretty much set it up.

OpenStudy (jhannybean):

I'm not really sure where to go with this problem now..

OpenStudy (anonymous):

Is that really the curl? The tangent vector is \(\langle f_x, -1, f_z\rangle\) then?

OpenStudy (jhannybean):

This example problem is telling me I need to use the cross product to figure out my vector, yes, is that what you are referring to?

OpenStudy (anonymous):

Yeah, I'm talking about the curl of the tangent vectors of the surface

OpenStudy (jhannybean):

I'm confused.

OpenStudy (anonymous):

I don't quite get where you are stuck?

OpenStudy (anonymous):

\[ f(x,z) = 4x+z^2 \]

OpenStudy (jhannybean):

Oh now I see where \(\langle f_ x, -1, f_z\rangle\) comes into the picture.

OpenStudy (jhannybean):

It's because i'm given a function in terms of the component \(y \hat j\)?

OpenStudy (anonymous):

Yeah, when we have a parametrization where two of the variables are in our x,y,z coordinate system, it end up having that normal vector.

OpenStudy (anonymous):

It's sort of special benefit you get from using the same coordinate system.

OpenStudy (jhannybean):

I see. Makes sense. I'll try working it out.

OpenStudy (jhannybean):

Do I need to start by taking the cross product?

OpenStudy (anonymous):

Have you done surface integrals before? Should I start from the beginning?

OpenStudy (jhannybean):

Yeah.. I was just working through the homework trying to understand as I went along.

OpenStudy (anonymous):

The same way that a path integral over the function \(f(x,y,z) = 1\) gives arc length, a surface integral over this function will give us surface area. Surface integrals are basically a special notation.

OpenStudy (anonymous):

\[ \iint_Sf ~dS \]Is a special notation. \(S\) stands for some surface we integrate over. \(f\) is the scalar function we are integrating over. Lastly \(dS\) is the change in surface area. Since we want surface area we set \(f=1\).

OpenStudy (anonymous):

First thing we will do is parametrize \(S\). We need two free variables to parametrize a surface. In general we will find a parametrization like \(\Phi(u,v) = \langle x(u,v), y(u,v), z(u,v)\rangle\). Where we bind the parametrization variables with: \(u_1 < u < u_2, v_1<v<v_2\).

OpenStudy (anonymous):

And finally \(\large dS = \left|\frac{\partial \Phi}{\partial u}\times \frac{\partial \Phi}{\partial v}\right|du~dv\)

OpenStudy (jhannybean):

Oh I see. It's becoming a little bit more clear now.

OpenStudy (anonymous):

\[ \iint_Sf ~dS = \int_{u_1}^{u_2}\int_{v_1}^{v_2}f~ \left|\frac{\partial \Phi}{\partial u}\times \frac{\partial \Phi}{\partial v}\right|du~dv \]

OpenStudy (anonymous):

Now, \(f\) is a function of \(x,y,z\), and we compose it with \(\Phi\) to get a function of \(u,v\)

OpenStudy (jhannybean):

Alright.

OpenStudy (anonymous):

In our case, we get surface area so \(f=1\). Also when we parametrize such that only one variable \(x,y,z\) is a function of the other two, we already know what the normal vector will be

OpenStudy (anonymous):

We know it would be \(\langle f_x,f_y,-1 \rangle\)

OpenStudy (anonymous):

That is if we had \(z=f(x,y)\). The variable that becomes a function of the other gets turned to \(-1\), the others are replaced with their partial derivatives of the parametrization function

OpenStudy (jhannybean):

Ahhh, yess.

OpenStudy (anonymous):

\[ |\langle f_x,f_y,-1\rangle| = \sqrt{\langle f_x,f_y,-1\rangle\cdot \langle f_x,f_y,-1\rangle} = \sqrt{f_x^2+f_y^2+1} \]

OpenStudy (anonymous):

Well, you can easy see this if you actually try to find the curl yourself

OpenStudy (anonymous):

\[ \begin{vmatrix} \mathbf i&\mathbf j&\mathbf k\\ \frac{\partial}{\partial x}x &\frac{\partial}{\partial x}y&\frac{\partial}{\partial x}f(x,y)\\ \frac{\partial}{\partial y}x &\frac{\partial}{\partial y}y&\frac{\partial}{\partial y}f(x,y)\\ \end{vmatrix} = \begin{vmatrix} \mathbf i&\mathbf j&\mathbf k\\ 1 &0&f_x\\ 0 &1&f_y\\ \end{vmatrix} = \langle f_x,f_y,-1\rangle \]

OpenStudy (anonymous):

My usage of \(f\) here is not the function we are integrating over, it's the parametrization: \[ \Phi (x,y) = \langle x,y,f(x,y)\rangle \]

OpenStudy (anonymous):

So going back to the actual problem at hand...

OpenStudy (anonymous):

\[ \Phi(x,y)=\langle x ,4x+z^2 , z\rangle \]

OpenStudy (jhannybean):

Oh, I was rereading your posts on how the variable that becomes the function turns into -1, and how you got z = -1. I understand that now.

OpenStudy (anonymous):

And \[ dS = \sqrt{\Bigg(\frac{\partial }{\partial x}\bigg(4x+z^2\bigg)\Bigg)^2+1+\Bigg(\frac{\partial }{\partial z}\bigg(4x+z^2\bigg)\Bigg)^2}dx~dz \]

OpenStudy (anonymous):

All together:\[ S=\iint_SdS = \int_0^1\int_0^1\sqrt{\Bigg(\frac{\partial }{\partial x}\bigg(4x+z^2\bigg)\Bigg)^2+1+\Bigg(\frac{\partial }{\partial z}\bigg(4x+z^2\bigg)\Bigg)^2}dx~dz \]Where \(S\) is the surface area of our parametrized surface

OpenStudy (jhannybean):

And you got the limits of integration because \(0 \le x \le 1\) and \(0\le z \le 1\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Basically, when you parametrize with the same coordinate system, you don't have to compute the curl of the tangent vectors because it will follow the pattern

OpenStudy (jhannybean):

Yeah, I see that now. And so if this equation was in terms of x, instead of y or z. We would have \(\langle -1, f_y, f_z\rangle\)

OpenStudy (anonymous):

But it's a special case. You shouldn't forget how you'd do a surface integral with a custom parametrization

OpenStudy (anonymous):

\[ \Phi(u,v) = \langle \cos u\sin v,\sin u\sin v,\cos v\rangle \]

OpenStudy (anonymous):

\[ \begin{vmatrix} \mathbf i&\mathbf j&\mathbf k\\ \frac{\partial}{\partial u}\cos u \sin v &\frac{\partial}{\partial u}\sin u \sin v&\frac{\partial}{\partial u}\cos v\\ \frac{\partial}{\partial v}\cos u \sin v &\frac{\partial}{\partial v}\sin u \sin v&\frac{\partial}{\partial v}\cos v\\ \end{vmatrix} \]

OpenStudy (jhannybean):

< sin(u)'sin(v) * (cos(v))' - (0)*sin(u)sin(v)' , ........>

OpenStudy (anonymous):

\[ \int_C ds \]

OpenStudy (anonymous):

\[ \int_C dr \]

OpenStudy (anonymous):

\[ \iint_S dS \]

OpenStudy (anonymous):

\[ \int_Cf~dr \]

OpenStudy (anonymous):

\[ \iint _Sf~dS \]

OpenStudy (anonymous):

\[ \int_C\mathbf F\cdot d\mathbf r \]

OpenStudy (anonymous):

\[ \iint_S\mathbf F\cdot d\mathbf S \]

OpenStudy (jhannybean):

\[\int_0^1 \sqrt{17+4z^2}dz\]

OpenStudy (jhannybean):

\[\int_0^1 \sqrt{4(\frac{17}{4} +z^2)}dz\]

OpenStudy (jhannybean):

\[2\int_0^1 \sqrt{\frac{17}{4} +z^2}dz\]

OpenStudy (anonymous):

\[ 1^2+\tan^2\theta = \sec^2\theta \]

OpenStudy (anonymous):

\[ \sqrt{17/4}^2+(\sqrt{17/4}\tan\theta)^2=(\sqrt{17/4}\sec\theta)^2 \]

OpenStudy (anonymous):

\[ \sin^2\theta+\cos^2=1 \]

OpenStudy (anonymous):

\[ \cos^2\theta=1-\sin^2\theta \]

OpenStudy (anonymous):

\[ \tan^2\theta+1=\sec^2\theta \]

OpenStudy (anonymous):

\[ 1-x^2\\ 1+x^2\\ x^2-1 \]

OpenStudy (anonymous):

\[ 1-x^2 \to 1-\sin^2\theta=\cos^2\theta\\ 1+x^2\to 1+\tan^2\theta=\sec^2\theta \\ x^2-1 \to \sec^2\theta-1=\tan^2\theta \]

OpenStudy (anonymous):

\[ z = \sqrt{17/4}\tan\theta \]

OpenStudy (anonymous):

\[ a^2-x^2 \to a^2-a^2\sin^2\theta=a^2\cos^2\theta\\ a^2+x^2\to a^2+a^2\tan^2\theta=a^2\sec^2\theta \\ x^2-a^2 \to a^2\sec^2\theta-a^2=a^2\tan^2\theta \]